# TabEBM: A Tabular Data Augmentation Method with Distinct Class-Specific Energy-Based Models

### v2

@@ Abstract
Data collection is often difficult in critical fields such as medicine, physics, and chemistry. As a result, classification methods usually perform poorly with these small datasets, leading to weak predictive performance. Increasing the training set with additional synthetic data, similar to data augmentation in images, is commonly believed to improve downstream classification performance. However, current tabular generative methods that learn either the joint distribution $p(\mathbf{x},y)$ or the class-conditional distribution $p(\mathbf{x}\mid y)$ often overfit on small datasets, resulting in poor-quality synthetic data, usually worsening classification performance compared to using real data alone. To solve these challenges, we introduce TabEBM, a novel class-conditional generative method using Energy-Based Models (EBMs). Unlike existing methods that use a shared model to approximate all class-conditional densities, our key innovation is to create distinct EBM generative models for each class, each modelling its class-specific data distribution individually. This approach creates robust energy landscapes, even in ambiguous class distributions. Our experiments show that TabEBM generates synthetic data with higher quality and better statistical fidelity than existing methods. When used for data augmentation, our synthetic data consistently improves the classification performance across diverse datasets of various sizes, especially small ones. Code is available at https://github.com/andreimargeloiu/TabEBM.
@@ 1 Introduction
Many scientific domains within medicine, physics, and chemistry often rely on intricate and challenging data acquisition procedures [5, 50, 4, 32, 76, 12], that typically render small-size tabular datasets [5, 46]. Using these to train machine learning models, that can aid in tasks such as disease diagnosis [52, 37], material property prediction [35], and chemical compound classification [11], can lead to poor performance [74, 52, 37]. In the case of learning tasks which leverage image and text data, a standard remedy to address performance issues due to data scarcity is employing data augmentation techniques [72, 73, 60, 71] that generate additional synthetic samples from existing data. However, applying data augmentation to tabular data introduces additional challenges, as tabular datasets are often very diverse and lack explicit symmetries [8]. Consequently, existing tabular data augmentation methods often yield mixed results and can even degrade model performance [51, 71, 48], hindering their widespread adoption.
@@
Tabular augmentation typically involves training generative models to approximate either the joint distribution $p({\mathbf{x}},y)$ [85, 24] or the class-conditional distribution $p({\mathbf{x}}|y)$ [85, 42, 83, 47, 48]. A key challenge of joint distribution methods is maintaining the original training label distribution, as sampling from such generators can produce label distributions that deviate from the original and even fail to generate data for specific classes (see Appendix C for an example). These issues compromise the effectiveness of data augmentation [51] by undermining the label accuracy and distribution. On the other hand, while class-conditional models learning $p({\mathbf{x}}|y)$ preserve the stratification of the original data, they often employ a shared model to represent all class-conditional densities. This, however, can lead to overfitting, particularly in imbalanced datasets where the model may prioritise more frequent classes [21], ignoring unique features needed for generating label-invariant samples. Additionally, in datasets with limited data, this can lead to mode collapse [68, 70], where the model does not effectively capture the diversity of each class [70], and thus tends to perform poorly in a multi-class setting.
@@
Figure 2: An overview of the proposed TabEBM. We learn distinct class-specific Energy-Based Models (EBMs) $E_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}blue}}}({\mathbf{x}})$ and $E_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{red}}}({\mathbf{x}})$ exclusively on the points of their respective class. Each EBM approximates a class-conditional distribution $p({\mathbf{x}}|y)$. TabEBM allows synthetic data generation by sampling from the estimated distributions for each class $p({\mathbf{x}}|y={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{blue}})$ and $p({\mathbf{x}}|y={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{red}})$.
@@ 2.1 Preliminaries on Energy-Based Models
An important observation is that energy-based models can utilise the same model architectures as standard classification models [29]. Typically, the logits $f_{\theta}({\mathbf{x}})[y]$ from a classification model define a discriminative distribution through the softmax function, expressed as $p_{\theta}(y|{\mathbf{x}})=\mathrm{softmax}(f_{\theta}({\mathbf{x}})[y])$. Intriguingly, these same logits can be reinterpreted to define an energy-based model for the joint distribution $p({\mathbf{x}},y)$. This is achieved by setting the energy function to $E({\mathbf{x}},y)=-f_{\theta}({\mathbf{x}})$. Furthermore, the energy function for the marginal distribution $p({\mathbf{x}})$ is obtained by marginalising over $p({\mathbf{x}},y)$, resulting in $E({\mathbf{x}})=-\text{LogSumExp}_{y^{\prime}}f_{\theta}({\mathbf{x}})[y^{\prime}]$.
@@ 2.2 Distinct Class-Specific Energy-Based Models
For the binary classifier $f^{c}_{\theta}(\cdot)$ in the surrogate binary classification, we use TabPFN [33], a pre-trained tabular in-context model. Note that TabPFN is intended for inference only, with no updates to its parameters (see Section 4 for more details about TabPFN). In this context, “training” the TabPFN classifier is analogous to the K Nearest Neighbour algorithm, which simply performs inference based on a training dataset provided to the model. We apply TabPFN multiple times on separate datasets $\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{C}\}$ to obtain multiple classifiers $\{f_{\theta}^{1},f_{\theta}^{2},\ldots,f_{\theta}^{C}\}$. In Section 3.4, we explore why reinterpreting TabPFN’s logits, trained on our surrogate binary tasks, can be useful for estimating an energy function. We emphasise that TabEBM is a general method, capable of using any gradient-based classifier that computes logits (using Equation 3), and is not limited to TabPFN.
@@ 3 Experiments
Datasets. We utilise eight open-source datasets from OpenML [7] – across five domains: Medicine, Chemistry, Engineering, Language and Economics. We note that TabPFN almost drains the small-size OpenML datasets for its meta-validation [33], which might lead to data leakage when evaluating TabEBM. To provide fair comparisons, we further select six leakage-free datasets from UCI [22]. These diverse datasets contain 7 to 77 features and 698 to 5500 samples across 2 to 26 classes. Five datasets contain both numerical and categorical features, while the remaining are numerical only. We further enlarge the evaluation scope by varying the degrees of data availability (i.e., $N_{\text{real}}$), leading up to 33 different test cases for the eight OpenML datasets. Section B.1 provide detailed descriptions.
@@
Data augmentation setup. Given $N_{\text{real}}$ real samples, we first train generators on the real training data and then generate $N_{\text{syn}}$ synthetic samples. For training the downstream predictors, we expand the real training split by adding the synthetic samples. The real validation data is used for early stopping, and the real test set is used for evaluating the predictor’s performance. The optimal $N_{\text{syn}}$ remains an open problem for tabular data [51, 71, 31]. Prior work [47, 48] mainly use synthetic sets with equivalent sizes to the real sets (i.e., $N_{\text{real}}=N_{\text{syn}}$). However, we observe that $N_{\text{real}}=N_{\text{syn}}$ can lead to highly unstable results, especially on small datasets we investigate. Recent work has used different $N_{\text{syn}}$ for various generators, such as by applying post-processing [31, 71]. In this work, we want to provide a head-to-head comparison of the effect of data augmentation across subsampled datasets of varying size $N_{\text{real}}\in\{20,50,100,200,500\}$. Therefore, we perform data augmentation with a large synthetic set ($N_{\text{syn}}=500$) across all splits, and the synthetic data has the same class distribution as the real training data. We provide an illustrative figure of the data splitting setup in Section B.2.
@@
3.1 Data Augmentation Improvement (Q1)
TabEBM effectively improves downstream performance across sample sizes, especially for very low-sample-size regimes. Table 1 and Figure 4 (Left) show that TabEBM exhibits competitive performance in data augmentation, generally achieving the highest downstream accuracy and average rank across most datasets and sample sizes. Notably, TabEBM is the only generator that consistently improves performance across sample sizes. A key observation is that most modern benchmark generators even underperform the Baseline, indicating poor approximated distributions in the low-sample-size regime. Moreover, TabEBM achieves the largest overall performance improvement on six leakage-free UCI datasets, further supporting its effectiveness (see Section D.5.2 for details).
@@ 3.1 Data Augmentation Improvement (Q1)
TabEBM effectively improves downstream performance across the number of classes, especially for more than ten classes. Figure 4 (Right) shows that TabEBM consistently outperforms the Baseline with notable improvements, particularly in datasets with more than ten classes. In contrast, an increased number of classes tends to cause a performance degradation in the benchmark generators.
@@ 3.2 Statistical Fidelity (Q2)
We evaluate the fidelity of synthetic data by focusing on two aspects of synthetic data: similarity to real train data and real test data. We evaluate the similarity between real and synthetic data via (i) average inverse of the Kullback–Leibler Divergence (inverse KL) [17], (ii) p-value of Kolmogorov–Smirnov test (KS test) [39] and (iii) p-value of Chi-squared test ($\chi^{2}$ test) [55]. The full results, including $\chi^{2}$ test and the similarity between real test data and synthetic data, are in Section D.6. For all three metrics, a bigger value denotes that synthetic data is more likely to have the same distribution as real data.
In Figure 7 (a1&a2), TabEBM consistently achieves the highest accuracy and distribution similarity between real train data and synthetic data, showing that TabEBM learns the distributions of real train data better than benchmark generators. In Section D.6, we further show that TabEBM remains the most competitive method in similarity between real test data and synthetic data. This indicates that TabEBM can extrapolate beyond real train data and thus generate synthetic data from a more general distribution that aligns with both train and test data. This extrapolation ability also explains why TabEBM can outperform Baseline via data augmentation (Section 3.1).
@@ 3.3 Privacy Preservation (Q3)
More broadly, data privacy is a critical concern for organisations and governments handling sensitive data [75]. Privacy-preserving synthetic data also allows researchers and practitioners to bypass ethical and logistical issues while enabling model training and testing [38]. We further explore the use of TabEBM-generated data for data sharing, where only synthetic data is accessible for downstream users [75, 89, 86, 23, 42]. In this case, downstream models are trained exclusively on synthetic data.
@@
Figure 7 (b1&b2) shows that TabEBM consistently finds a better trade-off between accuracy and privacy preservation. Notably, the “train-on-synthetic, test-on-real” scenario poses a greater challenge for generators in achieving high accuracy because real data is inaccessible for model training and data augmentation. Despite this difficulty, TabEBM is the only generator that surpasses the overall performance of training on real data (i.e., Baseline). The relatively high DCR for TabEBM indicates that it can extrapolate beyond real train data, aligning with the finding that TabEBM’s synthetic data is statistically similar to real test data (Section 3.2). These results further suggest that TabEBM learns the general distribution of real data, and TabEBM can generate high-quality synthetic data suitable for various purposes, including privacy preservation.
Figure 7: (a1&a2): Median inverse KL and KS test vs. mean normalised balanced accuracy improvement (%) between real train data and synthetic data. (b1&b2): Median DCR and $\delta$-presence vs. mean normalised balanced accuracy change (%) between real train data and synthetic data. Note that “accuracy improvement” is for data augmentation, and “accuracy change” is for data sharing. The figures with with standard deviations are in Section D.4. TabEBM generates high-fidelity synthetic data that can also be used for privacy preservation.
@@ 3.4 Why is TabEBM effective for estimating Energy-Based Models? (Q4)
We found it essential to place the negative samples far from the real data, since TabPFN, which pre-trained to approximate Bayesian inference [33], has its confidence influenced by the distance from the training data [53]. Figure 8 (left) shows that TabPFN outputs high logit values near the real data. As the distance from the real data increases, the logit $f({\mathbf{x}})[1]$ decreases smoothly until the two logits become similar, making the classifier uncertain (because the class probabilities become equal). The right figure shows that TabEBM’s inferred density drops significantly as the maximum logit decreases, because $p_{c}({\mathbf{x}})\propto(\exp(f({\mathbf{x}})[0])+\exp(f({\mathbf{x}})[1]))$ from Equation 3. Since SGLD sampling performs gradient ascent on the density, the TabEBM-generated samples will be close to the real data. These findings are consistent across datasets (see Section D.3), where TabPFN’s logits remain positive, with similar ranges and a relatively constant sum as distance increases, warranting further investigation. Overall, TabPFN’s distance-based uncertainty is useful for inferring accurate energy functions within our TabEBM framework. Since TabEBM can be paired with any other gradient-based classifier that produces logits, we leave these extensions for future work.

### v3

@@ Abstract
Data collection is often difficult in critical fields such as medicine, physics, and chemistry, yielding typically only small tabular datasets. However, classification methods tend to struggle with these small datasets, leading to poor predictive performance. Increasing the training set with additional synthetic data, similar to data augmentation in images, is commonly believed to improve downstream tabular classification performance. However, current tabular generative methods that learn either the joint distribution $p(\mathbf{x},y)$ or the class-conditional distribution $p(\mathbf{x}\mid y)$ often overfit on small datasets, resulting in poor-quality synthetic data, usually worsening classification performance compared to using real data alone. To solve these challenges, we introduce TabEBM, a novel class-conditional generative method using Energy-Based Models (EBMs). Unlike existing tabular methods that use a shared model to approximate all class-conditional densities, our key innovation is to create distinct EBM generative models for each class, each modelling its class-specific data distribution individually. This approach creates robust energy landscapes, even in ambiguous class distributions. Our experiments show that TabEBM generates synthetic data with higher quality and better statistical fidelity than existing methods. When used for data augmentation, our synthetic data consistently leads to improved classification performance across diverse datasets of various sizes, especially small ones. Code is available at https://github.com/andreimargeloiu/TabEBM.
@@ 1 Introduction
Many scientific domains within medicine, physics, and chemistry often rely on intricate and challenging data acquisition procedures [5, 50, 4, 32, 76, 12] that typically render small-size tabular datasets [5, 46]. Using these to train machine learning models that can aid in tasks such as disease diagnosis [52, 37], material property prediction [35], and chemical compound classification [11], can lead to poor performance [74, 52, 37]. In the case of learning tasks which leverage image and text data, a standard remedy to address performance issues due to data scarcity is employing data augmentation techniques [72, 73, 60, 71] that generate additional synthetic samples from existing data. However, applying data augmentation to tabular data introduces additional challenges, as tabular datasets are often very diverse and lack explicit symmetries [8], such as rotations or translations seen in images. Consequently, existing tabular data augmentation methods often yield mixed results and can even degrade model performance [51, 71, 48], hindering their widespread adoption.
@@
Tabular augmentation typically involves training generative models to approximate either the joint distribution $p({\mathbf{x}},y)$ [85, 24] or the class-conditional distribution $p({\mathbf{x}}|y)$ [85, 42, 83, 47, 48]. A key challenge of joint distribution methods is maintaining the original training label distribution, as sampling from such generators can produce label distributions that deviate from the original and even fail to generate data for specific classes (see Appendix C for an example). These issues compromise the effectiveness of data augmentation [51] by undermining the label accuracy and distribution. On the other hand, while class-conditional models that learn $p({\mathbf{x}}|y)$ preserve the stratification of the original data, they often employ a shared model to represent all class-conditional densities. This, however, can lead to overfitting, particularly in imbalanced datasets where the model may prioritise more frequent classes [21], ignoring unique features needed for generating label-invariant samples. Additionally, in datasets with limited data, this can lead to mode collapse [68, 70], where the model does not effectively capture the diversity of each class [70], and thus tends to perform poorly in a multi-class setting.
@@
Figure 2: An overview of TabEBM. We learn distinct class-specific Energy-Based Models (EBMs) $E_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}blue}}}({\mathbf{x}})$ and $E_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{red}}}({\mathbf{x}})$ exclusively on the points of their respective class. Each EBM approximates a class-conditional distribution $p({\mathbf{x}}|y)$. TabEBM allows synthetic data generation by sampling from the estimated distributions for each class $p({\mathbf{x}}|y={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{blue}})$ and $p({\mathbf{x}}|y={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{red}})$.
@@ 2.1 Preliminaries on Energy-Based Models
An important observation is that energy-based models can utilise the same model architectures as standard classification models [29]. Typically, the logits $f_{\theta}({\mathbf{x}})[y]$ from a classification model define a discriminative distribution through the softmax function, expressed as $p_{\theta}(y|{\mathbf{x}})=\mathrm{softmax}(f_{\theta}({\mathbf{x}})[y])$. Intriguingly, these same logits can be reinterpreted to define an energy-based model for the joint distribution $p({\mathbf{x}},y)$. This is achieved by setting the energy function to $E({\mathbf{x}},y)=-f_{\theta}({\mathbf{x}})$. Furthermore, the energy function for the marginal distribution $p({\mathbf{x}})$ is obtained by marginalising over $p({\mathbf{x}},y)$, resulting in $E({\mathbf{x}})=-\mathrm{{LogSumExp}}_{y^{\prime}}f_{\theta}({\mathbf{x}})[y^{\prime}]$.
@@ 2.2 Distinct Class-Specific Energy-Based Models
For the binary classifier $f^{c}_{\theta}(\cdot)$ in the surrogate binary classification, we use TabPFN [33], a pre-trained tabular in-context model. Note that TabPFN is intended for inference only, with no updates to its parameters (see Section 4 for more details about TabPFN). In this context, “training” the TabPFN classifier is analogous to the K-Nearest Neighbour algorithm, which simply performs inference based on a training dataset provided to the model. We apply TabPFN multiple times on separate datasets $\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{C}\}$ to obtain multiple classifiers $\{f_{\theta}^{1},f_{\theta}^{2},\ldots,f_{\theta}^{C}\}$. In Section 3.4, we explore why reinterpreting TabPFN’s logits, trained on our surrogate binary tasks, can be useful for estimating an energy function. We emphasise that TabEBM is a general method, capable of using any gradient-based classifier that computes logits (using Equation 3), and is not limited to TabPFN.
@@ 3 Experiments
Datasets. We utilise eight open-source tabular datasets from OpenML [7] across five domains: Medicine, Chemistry, Engineering, Language and Economics. As TabPFN utilises many small-size OpenML datasets in its meta-validation [33], it can lead to data leakage when evaluating TabEBM. Therefore, to provide fair comparisons, we select six additional leakage-free datasets from UCI [22]. These diverse datasets contain 7 to 77 features and 698 to 5500 samples across 2 to 26 classes. Five datasets contain both numerical and categorical features, while the remaining are numerical only. We further enlarge the evaluation scope by varying the degrees of data availability (i.e., $N_{\text{real}}$), leading up to 33 different test cases for the eight OpenML datasets. Section B.1 provides detailed descriptions.
@@
Data augmentation setup. Given $N_{\text{real}}$ real samples, we first train generators on the real training data and then generate $N_{\text{syn}}$ synthetic samples. For training the downstream predictors, we expand the real training split by adding the synthetic samples. The real validation data is used for early stopping, and the real test set is used for evaluating the predictor’s performance. The optimal $N_{\text{syn}}$ remains an open problem for tabular data [51, 71, 31]. Prior works [47, 48] mainly use synthetic sets with equivalent sizes to the real sets (i.e., $N_{\text{real}}=N_{\text{syn}}$). However, we observe that $N_{\text{real}}=N_{\text{syn}}$ can lead to highly unstable results, especially on small datasets that we investigate. Recent work has used different $N_{\text{syn}}$ for various generators, such as by applying post-processing [31, 71]. In this work, we want to provide a head-to-head comparison of the effect of data augmentation across subsampled datasets of varying sizes $N_{\text{real}}\in\{20,50,100,200,500\}$. Therefore, we perform data augmentation with a large synthetic set ($N_{\text{syn}}=500$) across all splits, and the synthetic data has the same class distribution as the real training data. We provide an illustrative figure of the data splitting setup in Section B.2.
3.1 Data Augmentation Improvement (Q1)
TabEBM effectively improves downstream performance across sample sizes, especially for very low-sample-size regimes. Table 1 and Figure 4 (Left) show that TabEBM exhibits competitive performance in data augmentation, generally achieving the highest downstream accuracy and average rank across most datasets and sample sizes. Notably, TabEBM is the only generator that consistently improves performance across sample sizes. A key observation is that most modern benchmark generators underperform even the Baseline, indicating poor approximated distributions in the low-sample-size regime. Moreover, TabEBM achieves the largest overall performance improvement on six leakage-free UCI datasets, further supporting its effectiveness (see Section D.5.2 for details).
@@ 3.1 Data Augmentation Improvement (Q1)
TabEBM effectively improves downstream performance across any number of classes, especially for more than ten classes. Figure 4 (Right) shows that TabEBM consistently outperforms the Baseline with notable improvements, particularly in datasets with more than ten classes. In contrast, an increased number of classes tends to cause a performance degradation in the benchmark generators.
@@ 3.2 Statistical Fidelity (Q2)
We evaluate the fidelity of synthetic data by measuring the similarity of synthetic data to real train data and to real test data (Figure 7). We evaluate this similarity via (i) average inverse of the Kullback–Leibler Divergence (inverse KL) [17], (ii) p-value of Kolmogorov–Smirnov test (KS test) [39] and (iii) p-value of Chi-squared test ($\chi^{2}$ test) [55]. For full numerical results, including $\chi^{2}$ test, see Section D.6. For all three metrics, a bigger value denotes that synthetic data is more likely to have the same distribution as real data.
In Figure 7 (a1&a2), TabEBM consistently exhibits the highest accuracy and distribution similarity between real train data and synthetic data, indicating that TabEBM learns the distributions of real train data better than benchmark generators. In Section D.6, we further show that TabEBM remains the most competitive method in similarity between real test data and synthetic data. This indicates that TabEBM can extrapolate beyond real train data and thus generate synthetic data from a more general distribution that aligns with both train and test data. This extrapolation ability also explains why TabEBM can outperform Baseline via data augmentation (Section 3.1).
@@ 3.3 Privacy Preservation (Q3)
More broadly, data privacy is a critical concern for organisations and governments handling sensitive data [75]. Privacy-preserving synthetic data allows researchers and practitioners to bypass ethical and logistical issues while enabling model training and testing [38]. We further explore the use of TabEBM-generated data for data sharing, where only synthetic data is accessible for downstream users [75, 89, 86, 23, 42]. In this case, downstream models are trained exclusively on synthetic data.
@@
Figure 7 (b1&b2) shows that TabEBM consistently finds a better trade-off between accuracy and privacy preservation. Notably, the “train-on-synthetic, test-on-real” scenario poses a greater challenge for generators in achieving high accuracy because real data is inaccessible for model training and data augmentation. Despite this difficulty, TabEBM is the only generator that surpasses the overall performance of training on real data (i.e., Baseline). The relatively high DCR for TabEBM indicates that it can extrapolate beyond real train data, aligning with the finding that TabEBM’s synthetic data is statistically similar to real test data (Section 3.2). These results further suggest that TabEBM learns the general distribution of real data, and can generate high-quality synthetic data suitable for various purposes, including privacy preservation.
Figure 7: (a1&a2): Median inverse KL and KS test vs. mean normalised balanced accuracy improvement (%) between real train data and synthetic data. (b1&b2): Median DCR and $\delta$-presence vs. mean normalised balanced accuracy change (%) between real train data and synthetic data. Note that “accuracy improvement” is for data augmentation, and “accuracy change” is for data sharing. Complete results with standard deviations are in Section D.4. TabEBM generates high-fidelity synthetic data that can also be used for privacy preservation.
@@ 3.4 Why is TabEBM effective for estimating Energy-Based Models? (Q4)
We found it essential to place the negative samples far from the real data, since TabPFN, which is pre-trained to approximate Bayesian inference [33], has its confidence influenced by the distance from the training data [53]. Figure 8 (left) shows that TabPFN outputs high logit values near the real data. As the distance from the real data increases, the logit $f({\mathbf{x}})[1]$ decreases smoothly until the two logits become similar, making the classifier uncertain (because the class probabilities become equal). Figure 8 (right) shows that TabEBM’s inferred density drops significantly as the maximum logit decreases, because $p_{c}({\mathbf{x}})\propto(\exp(f({\mathbf{x}})[0])+\exp(f({\mathbf{x}})[1]))$ from Equation 3. Since SGLD sampling performs gradient ascent on the density, the TabEBM-generated samples will be close to the real data. These findings are consistent across datasets (see Section D.3), where TabPFN’s logits remain positive, with similar ranges and a relatively constant sum as distance increases, warranting further investigation. Overall, TabPFN’s distance-based uncertainty is useful for inferring accurate energy functions within our TabEBM framework. Since TabEBM can be paired with any other gradient-based classifier that produces logits, we leave these extensions for future work.